prettygirl09:

halp math e.e p=2L+2w L=20 w=14

Joe348:

I WANNA DO ITTT

xxDeppressionxx:

Simplify 2×20 to 40. p=40+2×14 Simplify 2×14 to 28. p=40+28 Simplify 40+28 to 68. p=68

Joe348:

\[2(20)+2(14)\]

prettygirl09:

@xxdeppressionxx wrote:
Simplify 2×20 to 40. p=40+2×14 Simplify 2×14 to 28. p=40+28 Simplify 40+28 to 68. p=68
yay thankssss

xxDeppressionxx:

╭━╮╱╭╮╱╱╱╱╱╱╱╱╱╱╱╭╮╱╭╮ ┃┃╰╮┃┃╱╱╱╱╱╱╱╱╱╱╱┃┃╱┃┃ ┃╭╮╰╯┣━━╮╭━━┳━┳━━┫╰━┫┃╭━━┳╮╭╮ ┃┃╰╮┃┃╭╮┃┃╭╮┃╭┫╭╮┃╭╮┃┃┃┃━┫╰╯┃ ┃┃╱┃┃┃╰╯┃┃╰╯┃┃┃╰╯┃╰╯┃╰┫┃━┫┃┃┃ ╰╯╱╰━┻━━╯┃╭━┻╯╰━━┻━━┻━┻━━┻┻┻╯ ╱╱╱╱╱╱╱╱╱┃┃ ╱╱╱╱╱╱╱╱╱╰╯

AnishGC:

Perimeter = 2L +2W Perimeter = 2(20) + 2(14) Perimeter = 40 + 28 Perimeter = 68 units